Perdon Corporation manufactures safes—large mobile safes, and large walk-in stationary bank safes. As part of its annual budgeting process, Perdon is analyzing the profitability of its two products. Part of this analysis involves estimating the amount of overhead to be allocated to each product line. The information shown below relates to overhead.

Mobile Safes

Walk-in Safes

Units planned for production 200 50
Material moves per product line 300 200
Purchase orders per product line 450 350
Direct labor hours per product line 800 1,700

(a) The total estimated manufacturing overhead was $260,000. Under traditional costing (which assigns overhead on the basis of direct-labor hours), what amount of manufacturing overhead costs are assigned to: (Round intermediate calculations and answers to 2 decimal places, e.g. $12.25.)

(1) One mobile safe PER UNIT
(2) One walk-in safe PER UNIT

(b) The total estimated manufacturing overhead of $260,000 was comprised of $160,000 for material-handling costs and $100,000 for purchasing activity costs. Under activitybased costing (ABC): (Round intermediate calculations and answers to 2 decimal places, e.g. $12.25.)

(1) What amount of material handling costs are assigned to:

(a) One mobile safe
(b) One walk-in safe

(2) What amount of purchasing activity costs are assigned to:

(a) One mobile safe
(b) One walk-in safe
Answer #1

Solution :

(a) Traditional Costing :

Total Estimated OH     = $ 260,000

Overhead Allocated = (Total OH / Total Direct Labor Hours) * Labor Hours in Product

One Mobile Safe         = [$ 260,000 / (800 + 1,700)] * 800 = $ 83,200

Walk in Safe              = [$ 260,000 / (800 + 1,700)] * 1,700 = $ 176,800

(1) OH Allocated to One Mobile Safe = Total OH Allocated to Mobile Safe / No. of Unit of One Mobile Safe

                                                      = $ 83,200 / 200

= $ 416 per Unit

(2) OH Allocated to One Mobile Safe = Total OH Allocated to Mobile Safe / No. of Unit of Walk in Mobile Safe

                                                      = $ 176,800 / 50

= $ 3,536 per Unit

(b) ABC Costing :

(1) Total Material Handling Estimated OH     = $ 160,000

Overhead Allocated = (Total Material Handling OH / Total Material Moves) * Material Moves in Product

One Mobile Safe         = [$ 160,000 / (300 + 200)] * 300 = $ 96,000

Walk in Safe              = [$ 160,000 / (300 + 200)] * 200 = $ 64,000

(a) OH Allocated to One Mobile Safe = Material Handling OH Allocated to Mobile Safe / No. of Unit of One Mobile Safe

                                                      = $ 96,000 / 200

= $ 480 per Unit

(b) OH Allocated to One Mobile Safe = Material Handling OH Allocated to Mobile Safe / No. of Unit of Walk in Mobile Safe

                                                      = $ 64,000 / 50

= $ 1,280 per Unit

(2) Total Purchasing Estimated OH     = $ 100,000

Overhead Allocated = (Total Purchasing OH / Total Purchases Orders) * Purchases orders in Product

One Mobile Safe         = [$ 100,000 / (450 + 350)] * 450 = $ 56,250

Walk in Safe              = [$ 100,000 / (450 + 350)] * 350 = $ 43,750

(a) OH Allocated to One Mobile Safe = Purchasing OH Allocated to Mobile Safe / No. of Unit of One Mobile Safe

                                                      = $ 56,250 / 200

= $ 281.25 per Unit

(b) OH Allocated to One Mobile Safe = Purchasing OH Allocated to Mobile Safe / No. of Unit of Walk in Mobile Safe

                                                      = $ 43,750 / 50

= $ 875 per Unit

Please vote up and write your doubts in comment section. Would be glad to help you further.